OpenStudy (anonymous):

CHECK MY WORK!! The length of a rectangle is 24 units. Can the perimeter x of the rectangle be 60 units when its width y is 11 units? No, the rectangle cannot have x = 60 and y = 11 because x = 48 + 2y No, the rectangle cannot have x = 60 and y = 11 because x = 24 + 2y Yes, the rectangle can have x = 60 and y = 11 because x + y is greater than 48 Yes, the rectangle can have x = 60 and y = 11 because x + y is greater than 24 I know that it is a or b, right?

jimthompson5910 (jim_thompson5910):

hint: perimeter (P) P = 2*L + 2*W L = length W = width

OpenStudy (anonymous):

a??

OpenStudy (anonymous):

@jim_thompson5910 a?

jimthompson5910 (jim_thompson5910):

yes P = x L = 24 W = y P = 2*L + 2*W x = 2*24 + 2*y x = 48 + 2y

jimthompson5910 (jim_thompson5910):

you'll find that if L = 24 and W = 11, then the perimeter is actually 70 (not 60)

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

yw